10.0 of unknown metal was completely reacted with chlorine gas to yield(1.81x10^1) grams of the metal chloride. According to the law of conservation of mass, how many milliliters of chlorine reacted? Density of chlorine is 3.214g/L.
1
Expert's answer
2020-09-14T08:24:19-0400

Assuming you started with 10g of the metal;

Mass of the product is 18.1g,

Therefore mass of chlorine added is 18.1 -10 =8.1g


Volume of chlorine"=\dfrac{mass }{Density}" "=\dfrac{8.1g}{3.214g/L}" "=2.52L"